# Depth-Weighted Detection of Behaviours of Risk in People with Dementia using Cameras

### v1

@@ Abstract
The behavioural and psychological symptoms of dementia, such as agitation and aggression, present a significant health and safety risk in residential care settings. Many care facilities have video cameras in place for digital monitoring of public spaces, which can be leveraged to develop an automated behaviours of risk detection system that can alert the staff to enable timely intervention and prevent the situation from escalating. However, one of the challenges in our previous study was the presence of false alarms due to obstruction of view by activities happening close to the camera. To address this issue, we proposed a novel depth-weighted loss function to train a customized convolutional autoencoder to enforce equivalent importance to the events happening both near and far from the cameras; thus, helping to reduce false alarms and making the method more suitable for real-world deployment. The proposed method was trained using data from nine participants with dementia across three cameras situated in a specialized dementia unit and achieved an area under the curve of receiver operating characteristic of $0.852$, $0.81$ and $0.768$ for the three cameras. Ablation analysis was conducted for the individual components of the proposed method and the performance of the proposed method was investigated for participant-specific and sex-specific behaviours of risk detection. The proposed method performed reasonably well in detecting behaviours of risk in people with dementia motivating further research toward the development of a behaviours of risk detection system suitable for deployment in video surveillance systems in care facilities.
@@ 1 Introduction
Behaviours of risk detection is a challenging problem due to their diversity and rare occurrence, making supervised classification approaches inappropriate. An anomaly detection approach is more plausible, where a method can be trained to learn the normal behaviour characteristics and identify high variation cases as anomalies during testing [3]. In our previous work [4], a spatio-temporal autoencoder was used to detect behaviours of risk as anomalies in a single dementia participant and a single camera. However, there were significant issues with false positives (or false alarms), that could cause usability issues. In general, video anomaly detection (VAD) autoencoder approaches can become sensitive to the background, leading to high reconstruction errors when a major part of the background is blocked [5]. The objects or people closer to the camera cover a larger portion of the background than those further away, leading to significantly higher reconstruction errors and causing increased false positives. To tackle this problem, we propose a novel loss function where the depth of the pixels is used as weight to diminish the overshadowing effect caused by the larger area of the people/objects close to camera. The aim is to reduce false positives by enforcing equivalent importance in analyzing the events happening in the unit irrespective of the distance to camera. Further, in the absence of anomalies in the training set, we propose to utilize unusual activities that are not behaviours of risk, such as large objects or crowded scenes as proxy outliers to determine a threshold for detecting behaviours of risk. The performance for detecting behaviours of risk can vary across individuals depending upon the type of behaviours of risk, number of episodes recorded and demographic variables such as sex. To address this variability, we also conduct participant-specific and sex-specific analyses of the effectiveness of detecting behaviours of risk by the proposed method. The main contributions of this work are:
@@
    We propose a novel depth-weighted loss function, where pixel depth is used to enforce equivalent importance to both near and far events and reduce the false positives.
@@
    We use the outlier activities in the training data to determine a threshold. We expand our analysis to nine participants with dementia and three different cameras.
@@
    We conduct an ablation analysis to investigate the effectiveness of individual components of the proposed method in detecting behaviours of risk in PwD.
@@
    We perform experiments to analyze the performance of the proposed method for participant-specific and sex-specific behaviours of risk detection in PwD.
@@ 2 Related Work
The existing methods tried to reduce false positives by leveraging algorithmic designs or choosing the threshold depending on a desired false positive rate. However, as described in Section 1, the obstruction caused depending on the distance of objects/people from the camera can be a determinant factor in the increased false positives in VAD tasks. Hence, we utilize the depth of pixels to handle this issue and enforce equivalent importance to events happening far and close to the camera, thus reducing the false positives.
@@ 3.1 Data Collection
The video data used in this work was collected as part of a larger multimodal sensor study [6] at the Specialized Dementia Unit (SDU) at Toronto Rehabilitation Institute, University Health Network (UHN), Toronto, Canada, between November $2017$ and October $2019$. The participants were admitted to the SDU from long-term care homes for the evaluation and management of behavioural and psychological symptoms in advanced dementia. Inclusion criteria encompassed individuals aged over 55 years, diagnosed with dementia and with a history of agitation or aggressive behaviours. The participants who expressed dissent toward wearing the sensor device were excluded from the study. Data collection ceased under either of the following circumstances: absence of documented agitation for one week, repeated removal of the wristband, sickness, death or discharge from the unit, or upon completion of two months of data collection. The data collection was approved by the research ethics board (UHN REB#14-8483). Informed consent was obtained from substitute decision-makers for all participants. The staff also provided written consent for video recording within the unit. Additionally, all staff and decision-makers of participants shown in the images in the paper have granted written consent for the publication of their images.
@@
Figure 1: Different camera views.
@@ 3.2 Selected Dataset
For this study, we selected a subset from the larger video dataset comprising the top three cameras (see Fig. 1) based on the maximum documented behaviours of risk events. Table 1 presents the demographic information of the participants. The selected video data captured the behaviours of risk events for nine dementia participants with an age range of 6693669366–9366 – 93 years and a mean age of $81.22\pm 8.12$ years. There were three male and six female participants.
@@ 3.3 Data Annotation
Two research assistants independently reviewed the selected videos 30 minutes before and after the documented behaviours of risk events and annotated all the events corresponding to motor or physical agitation. Due to the lack of audio during data collection, it was not possible to annotate verbal aggression. Anvil [16] tool was used for the annotation of camera videos. Both the research assistants annotated the videos separately for the presence of behaviours of risk events and other activities, namely, the presence of unusual activities, crowded scenes and large objects. The two research assistants individually annotated $305$ and $383$ behaviours of risk events across three cameras and nine participants. For the final labelling, an event was labelled as behaviour of risk if either of the research assistants annotated it as a behaviour of risk. For cases where both the assistants annotated an event as behaviour of risk but disagreed in the start/end times of the annotation, the duration with common agreement was labelled as a behaviour of risk. An inter-rater agreement analysis was performed, where Cohen’s kappa [17], Krippendorff alpha [18] and percentage agreement metrics were calculated over the annotations of both research assistants. This was done to measure the agreeability of the two annotations over the categorization of the events into normal and behaviours of risk events. The Cohen’s kappa, Krippendorff alpha and percentage agreement were determined as $0.48$, $0.48$ and $0.99$, respectively. The value of Cohen’s kappa and Krippendorff alpha can range from $-1$ to $1$, where $1$ and $-1$ represent a perfect agreement and disagreement, and 00 represents the amount of agreement expected from random chance. A score between $0.41$ to $0.60$ for Cohen’s kappa is considered as moderate [17].
@@ 3.4 Data Preprocessing
The frames were sampled from the videos at $15$ frames per second, converted to grayscale, normalized to range $[0,1]$ (pixel values divided by $255$) and resized to $64\times 64$ resolution – this was done to minimize the computational expenses associated with trainable parameters. The frames were then stacked separately to form non-overlapping $5$-second windows ($75$ frames per window) (based on previous work [11]). The labels for behaviours of risk were refined in accordance with the $5$-second windows. Training and test sets were generated individually for the three cameras. The training sets were composed of normal behaviours from daily activities only. The test sets were composed of both normal and behaviours of risk events. The training set size was fixed around an average of $1224.41$ minutes across three cameras. The test set consists of a combination of behaviours of risk events from nine participants across three cameras along with normal behaviours. In the test sets, the available duration of annotated behaviours of risk events was limited. Hence, the duration of normal behaviour in test sets was chosen to keep a similar imbalance ratio (as in previous work [4]), which is approximately $7\%$ for behaviours of risk events. Table 2 shows the size of training and test sets for each of the three cameras. The number of $5$-second windows and types of behaviours of risk across participants and cameras in the test set are presented in Table 3.
@@
Participant# Sex Cam1 Cam2 Cam3 Type of behaviour of risk

### v2

@@ Abstract
The behavioural and psychological symptoms of dementia, such as agitation and aggression, present a significant health and safety risk in residential care settings. Many care facilities have video cameras in place for digital monitoring of public spaces, which can be leveraged to develop an automated behaviours of risk detection system that can alert the staff to enable timely intervention and prevent the situation from escalating. However, one of the challenges in our previous study was the presence of false alarms due to disparate importance of events based on distance. To address this issue, we proposed a novel depth-weighted loss to enforce equivalent importance to the events happening both near and far from the cameras; thus, helping to reduce false alarms. We further propose to utilize the training outliers to determine the anomaly threshold. The data from nine dementia participants across three cameras in a specialized dementia unit were used for training. The proposed approach obtained the best area under receiver operating characteristic curve performance of $0.852$, $0.81$ and $0.768$, respectively, for the three cameras. Ablation analysis was conducted for the individual components of the proposed approach and effect of frame size and frame rate. The performance of the proposed approach was investigated for cross-camera, participant-specific and sex-specific behaviours of risk detection. The proposed approach performed reasonably well in reducing false alarms. This motivates further research to make the system more suitable for deployment in care facilities.
@@ 1 Introduction
Behaviours of risk detection is a challenging problem due to their diversity and rare occurrence, making supervised classification approaches inappropriate. An anomaly detection approach is more plausible, where a method can be trained to learn the normal behaviour characteristics and identify high variation cases as anomalies during testing [3]. In our previous work [4], a spatio-temporal autoencoder was used to detect behaviours of risk as anomalies. However, there were significant issues with false positives (or false alarms), that could cause usability issues. In general, video anomaly detection (VAD) autoencoder approaches can become sensitive to the background, leading to high reconstruction errors when a major part of the background is blocked [5]. The objects or people closer to the camera cover a larger portion of the background than those further away, leading to significantly higher reconstruction errors and causing increased false positives. To tackle this problem, we propose a novel loss where the depth of the pixels is used as weight to diminish the overshadowing effect caused by the larger area of the people/objects close to camera. The aim is to reduce false positives by enforcing equivalent importance in analyzing the events happening in the unit irrespective of the distance to camera. In the absence of anomalies in the training set, we propose to utilize unusual activities that are not behaviours of risk, such as large objects or crowded scenes as proxy outliers to determine a threshold for detecting behaviours of risk. Expanding on our previous work [4] that analyzed behaviours of risk in one dementia participant using a single camera, here we extend the analysis to nine participants and three cameras. The performance for detecting behaviours of risk can vary across individuals depending upon the type of behaviours of risk, number of episodes recorded and demographic variables such as sex. To address this variability, we also conduct participant-specific and sex-specific analyses of the effectiveness of detecting behaviours of risk by the proposed approach. The main contributions of this work are:
@@
    We propose an approach using a novel depth-weighted loss to enforce equivalent importance to events irrespective of distance from camera and utilize the training outliers to determine the anomaly threshold.
@@
    We expand our analysis to nine participants with dementia and three different cameras.
@@
    We integrate the proposed approach into three different existing anomaly detection methods, referred to as their depth variants. The effectiveness of the proposed approach is then evaluated by comparing the performance of the depth variants with the existing methods.
@@
    We conduct an ablation analysis to investigate the effectiveness of individual components of the proposed approach and effect of frame size and frame rate in detecting behaviours of risk in PwD.
  5. 5.
    We analyze the performance of the proposed approach for generalization across cameras, participant-specific and sex-specific behaviours of risk detection in PwD.
@@ 2 Related Work
Video Anomaly Detection: VAD is a task in computer vision, aimed at identifying unusual or anomalous events in video sequences in the absence of anomalous samples during training. Among these, autoencoders and depth-based methods have emerged as prominent approaches. Liu et al. [12] proposed an appearance-motion united autoencoder framework for VAD which learned the appearance and motion normality jointly. Le and Kim [13] proposed an attention-based residual autoencoder which encoded both spatial and temporal information in a unified way. Schneider et al. [14] used foreground-aware loss and investigated how existing algorithms could leverage depth modality from Kinect camera to enhance detection performance. Denkovski et al. [15] proposed a multi-objective reconstruction and prediction-based loss to detect falls using multiple camera modalities including depth within a window of sequential frames.
@@
The existing methods tried to reduce false positives by leveraging algorithmic designs or choosing the threshold depending on a desired false positive rate. However, as described in Section 1, the sensitivity to the distance of objects/people from the camera can be a determinant factor in the increased false positives in VAD tasks, especially for scenes with large variations in depth, such as corridors or hallways. Hence, we utilize the depth of pixels to handle this issue and enforce equivalent importance to events happening far and close to the camera, thus reducing the false positives.
@@ 3.1 Data Collection
The video data used in this work was collected as part of a larger multimodal sensor study [6, 20] at the Specialized Dementia Unit (SDU) at Toronto Rehabilitation Institute, University Health Network (UHN), Toronto, Canada, between November $2017$ and October $2019$. The participants were admitted to the SDU from long-term care homes for the evaluation and management of behavioural and psychological symptoms in advanced dementia. Inclusion criteria encompassed individuals aged over 55 years, diagnosed with dementia and with a history of agitation or aggressive behaviours. The participants who expressed dissent to use of a wearable device (one of the study sensors, not part of this video-based study) were excluded from the study. Data collection ceased under either of the following circumstances: absence of documented agitation for one week, repeated removal of the wearable device, sickness, death or discharge from the unit, or upon completion of two months of data collection. The data collection was approved by the research ethics board (UHN REB#14-8483). Informed consent was obtained from substitute decision-makers for all participants. The staff also provided written consent for video recording within the unit. Additionally, all staff and decision-makers of participants shown in the images in the paper have granted written consent for the publication of their images.
@@
Figure 1: Three different installed camera views.
@@ 3.2 Selected Dataset
For this study, we selected a subset of videos, approximately $93.5$ hours for each camera, based on 30 minutes before and after each documented behaviour of risk event from the larger video dataset comprising the top three cameras (see Fig. 1). Table 1 presents the demographic information of the participants. The selected video data captured the behaviours of risk events for nine dementia participants with an age range of 6693669366–9366 – 93 years and a mean age of $81.22\pm 8.12$ years. There were three male and six female participants.
@@ 3.3 Data Annotation
Two research assistants were hired on a part-time basis and independently reviewed selected videos containing behaviour of risk events across Cam1, 2 and 3. We had the timestamps of each behaviour of risk event, so the annotators did not have to watch the entire video feed but only look at the selected segment for each behaviour of risk event. This activity took approximately two months and they annotated all the events corresponding to motor or physical agitation. Due to the lack of audio during data collection, it was not possible to annotate verbal aggression. Anvil [21] tool was used for the annotation of camera videos. Both the research assistants annotated the videos separately for the presence of behaviours of risk events and other activities, namely, the presence of unusual activities, crowded scenes and large objects. The two assistants individually annotated $305$ and $383$ behaviours of risk events across three cameras and nine participants. For the final labelling, an event was labelled as behaviour of risk if either of the research assistants annotated it as a behaviour of risk. For cases where both the assistants annotated an event as behaviour of risk but disagreed in the start/end times of the annotation, the duration with common agreement was labelled as a behaviour of risk. An inter-rater agreement analysis was performed, where Cohen’s kappa [22], Krippendorff alpha [23] and percentage agreement metrics were calculated over the annotations of both research assistants. The metrics were calculated based on the timestamps (in seconds) of the annotations of the two annotators. This was done to measure the agreeability of the two annotations over the categorization of the events into normal and behaviours of risk events. The Cohen’s kappa, Krippendorff alpha and percentage agreement were determined as $0.48$, $0.48$ and $0.99$, respectively. The value of Cohen’s kappa and Krippendorff alpha can range from $-1$ to $1$, where $1$ and $-1$ represent a perfect agreement and disagreement, and 00 represents agreement expected from random chance.
@@ 3.4 Data Preprocessing
The original resolution, frame rate and bit depth of the videos were $352\times 240$, $30$ frames per second (fps) and $24$, respectively. The frames were sampled from the videos at $15$ fps, selecting every second frame. The frames were then converted to grayscale, normalized to range $[0,1]$ (pixel values divided by $255$) and resized to $64\times 64$ resolution using the python programming language. This was done to minimize the computational expenses associated with trainable parameters. The frames were then stacked separately to form non-overlapping $5$-second windows ($75$ frames per window) (based on previous work [11]). The labels for behaviours of risk were refined in accordance with the $5$-second windows. Training and test sets were generated individually for the three cameras. The training sets were composed of normal behaviours from daily activities only. The test sets were composed of both normal and behaviours of risk events. The training set size was fixed around an average of $1224.41$ minutes across three cameras. The test set consists of a combination of behaviours of risk events from nine participants across three cameras along with normal behaviours. In the test sets, the available duration of annotated behaviours of risk events was limited. Hence, the duration of normal behaviour in test sets was chosen to keep a similar imbalance ratio (as in previous work [4]), which is approximately $7\%$ for behaviours of risk events. Table 2 shows the size of training and test sets for each of the three cameras. The test set was used to evaluate the performance of methods for detecting behaviours of risk. The number of $5$-second windows and types of behaviours of risk across participants and cameras in the test set are presented in Table 3.
@@
Participant-ID Sex Cam1 Cam2 Cam3 Type of behaviour of risk
@@
3.5 Behaviours of Risk Detection
In our previous work [11, 4], we used a customized spatio-temporal convolutional autoencoder (3DCAE) to detect behaviours of risk from data from one participant and a single camera. 3DCAE learned to reconstruct the input videos with normal behaviour during training. Hence, during testing, normal instances were expected to have a lower reconstruction error, while instances with high reconstruction error were treated as anomalies (behaviours of risk in our case). However, it showed high false positives due to sensitivity to activities such as crowded scenes, large objects, or other events happening near the camera due to reasons described in Section 1. Such instances are extremely common in normal day-to-day activities happening in care units. As such, there was a need to handle this issue to make the method suitable for real-world deployment. Hence in this work, we propose a novel depth-weighted loss, where the depth of the pixels is used as a weight factor to enforce equivalent importance to both behaviours of risk and non-behaviours of risk events happening far and near the camera. It is hypothesized that the weighted-depth loss will lead to a reduction in false positives in this problem. We also propose a new way to determine the threshold for detecting behaviours of risk as anomalies using the abnormal non-behaviours of risk activities in the training set as outliers. Hence, the proposed approach consists of the novel depth-weighted loss and utilizing the training outliers to determine the threshold.
Depth-weighted Anomaly Score: Existing autoencoder methods [4] employ reconstruction error to detect anomalous frames. This strategy inherently favours anomalies closer to the camera, as they occupy more pixels in the image plane. This bias poses a challenge, especially in scenes with highly varying distances from the camera, such as corridors (see Fig. 1). As a result, anomalies at different distances receive unequal anomaly scores, with those closer to the camera being unfairly favoured. To address this problem, we propose a loss where depth per pixel is used as weight during the calculation of reconstruction error to enforce equivalent importance to both near and far pixels.
[…]
[2pt]Refer to caption(a)
[…]
[2pt]Refer to caption(b)
Figure 2: (a) Illustration of the geometric relationship between a 3D prism (side length $L$) and the area of its projection in the image plane. (b) Two identical 3D prisms placed at different distances from camera position $C$ ($Z_{a}>Z_{b}$) produce a smaller projected area ($A_{a}<A_{b}$).
To understand the translation of an anomaly in $3$D space, we describe the relationship between the area $A$ depicting the anomaly in the image plane and its $3$D volume. For simplicity, we consider this $3$D volume as a square prism (see Fig. 2 (a)). Consider two of the vertex of the square prism represented as $x_{1}=(X_{1},Y_{1},Z_{1})$ and $x_{2}=(X_{2},Y_{2},Z_{2})$ in the $3$D space and their $2$D projections $u_{1}=(U_{1},V_{1})$ and $u_{2}=(U_{2},V_{2})$ in the image plane. In Fig. 2, $C$ refers to the camera position. The projection of a $3$D point $(X,Y,Z)$ onto a $2$D image plane is expressed by the image coordinates $(U,V)$ through the pinhole camera model [24] equations as,
$U=\frac{fX}{Z},\quad V=\frac{fY}{Z}$ (1)
where $f$ is the focal length and $Z$ is the distance from the $3$D object to the camera position. Assuming the difference in depth $Z$ between these vertices is negligible ($Z\approx Z_{1}\approx Z_{2}$), the projections on the image plane according to equation (1) are:
$U_{1}=\frac{fX_{1}}{Z},\quad U_{2}=\frac{fX_{2}}{Z}$ (2)
Given that the length between the vertices in the 3D space along the X-axis is $L=|X_{2}-X_{1}|$, the area $A$ of the square in the $2$D plane containing the anomaly is:
To analyze the sensitivity of the anomaly score $(AS)$ on the distance from the camera and therefore, on the area in the image plane, we can express the anomaly score of an area $A$ as,
$AS_{A}=\overline{AS}_{\text{pixel}}\cdot A$ (4)
where $\overline{AS}_{\text{pixel}}$ represents the average anomaly score per pixel, a constant value that depends solely on the anomaly itself, irrespective of its position in 3D space or its projection onto the image plane.
Consider an anomaly event occurring at different distances $Z_{a}$ and $Z_{b}$ from the camera. As it is the same event with same volume, it will have same value of $\overline{AS}_{\text{pixel}}$ and length $L$ at both distances. As illustrated in Fig. 2 (b), when projected in the image plane, this results in different area values $A_{a}$ and $A_{b}$ in the image plane, with $A_{a}<A_{b}$ if $Z_{a}>Z_{b}$. Since both anomalies have the same value of $\overline{AS}_{\text{pixel}}$, according to equation (4), the difference in area will result in a different anomaly score. To ensure depth invariance, we need the anomaly to have the same score $AS_{inv}$ regardless of its distance from the camera. This is achieved by introducing a factor $K$:
$AS_{inv}=K_{a}\cdot AS_{A_{a}}=K_{b}\cdot AS_{A_{b}}$ (5)
Combining the equations (3), (4) and (5), we can express $AS_{inv}$ as :
$AS_{inv}=K_{a}\overline{AS}_{\text{pixel}}\left(fL\right)^{2}\frac{1}{Z_{a}^{2}}=K_{b}\overline{AS}_{\text{pixel}}\left(fL\right)^{2}\frac{1}{Z_{b}^{2}}$ (6)
Therefore, given the constant value of $\overline{AS}_{\text{pixel}}$ and $\left(fL\right)^{2}$, to achieve depth invariance, the factors $K_{a}$ and $K_{b}$ must satisfy the following conditions:
$K_{a}=Z_{a}^{2},\quad K_{b}=Z_{b}^{2}$ (7)
Based on this, we propose the following novel loss where depth per pixel is used as weight during the calculation of reconstruction error as mean squared error to enable the methods to enforce equivalent importance to both near and far pixels.
$L_{dmse}=\frac{1}{N_{e}}\sum_{l=1}^{W}\sum_{i=1}^{S}\sum_{j=1}^{S}{\color[rgb]{0,0,0}\phi_{l,i,j}}\>\>(I_{l,i,j}-O_{l,i,j})^{2}$ (8)
where, $I$ is the input pixel, $O$ is the reconstructed pixel, $S$ is the spatial size, and $W$ represents the number of frames in a window. $N_{e}=W\times S\times S$ is the total number of pixels in a window. The depth-based weight hyperparameter $\phi$ can be expressed as a function of $Z^{2}$, i.e., $\phi=\mathfrak{f}(Z^{2})$. The value of $\phi$ is chosen using cross-validation from among different functions of $Z^{2}$, namely $\left(Z^{2},Z^{2}/2,\sqrt{Z^{2}}=Z\right)$, by leveraging the proxy validation set in a similar way as the threshold. The weighted-depth loss helped the method to treat both the near and far behaviours of risk events with similar importance, which in turn helped to reduce false positives. The dense prediction transformer [25] was used to estimate the depth map during both training and testing. During testing, the depth-weighted loss was used as anomaly score for detecting behaviours of risk.
Threshold Determination: In an anomaly detection approach, it can be challenging to determine an operating threshold due to the absence of anomalous samples in the training set. As discussed in Section 3.3, the data was annotated for other unusual activities, such as large objects, crowded scenes and other infrequent activities in the training set. Due to the unusual nature of these annotated activities relative to the activities of daily living in a dementia unit, we propose to treat these activities as proxy outliers to create a proxy validation set to determine the threshold. Following a similar approach as in our previous work [26], we chose each reconstruction error as a threshold during training and found the corresponding F1-score on the proxy validation set. Finally, the reconstruction error with the highest F1-score was chosen as the operating threshold. The operating threshold was then utilized during testing to detect behaviours of risk in the test set and the results were reported in Section 4.